# The Tautochrone of Huygens and Abel: From Constructive Geometry to Fractional Calculus

### v1

@@ 5 Concluding Remarks
Acknowledgments
L. R. E. gratefully acknowledges the hospitality of the Department of Molecular Sciences and Nanosystems at Ca’ Foscari University of Venice, Italy, and the Department of Materials, Environmental Sciences and Urban Planning (SIMAU) at the Polytechnic University of Marche (UnivPM), Ancona, Italy, where he served as a Visiting Scholar during the preparation of this work.
The research of F. M. has been carried out in the framework of the activities of the Italian National Group of Mathematical Physics (GNFM-INdAM).
Financial Support
@@ Acknowledgments
Data Availability Statement
@@
Generative Artificial Intelligence (AI)
This manuscript benefited from the use of Generative AI tools (ChatGPT) for translation assistance and grammar refinement, given that the authors are not native speakers of English.
Authors’ Biography
Luis Roberto Evangelista is Professor of Theoretical Physics at the PhD Preogram in Physics at the University of Maringá (UEM), Brazil. His research interest lie in complex fluids, complex systems, and history of physics, and include mathematical physics of liquid-crystals, diffusion problems, and adsorption-desorption phenomena.
His Orcid Number is 0000-0001-7202-1386
Francesco Mainardi is Professor (Retired) of Mathematical Physics at the University of Bologna, Italy, renowned for his pioneering work in fractional calculus. His research focuses on fractional differential equations, viscoelasticity, and anomalous diffusion. He played a key role in promoting the Mittag-Leffler and Wright functions in applied contexts. Mainardi authored influential books, including Fractional Calculus and Waves in Linear Viscoelasticity. He remains an active voice in the field, shaping theory and applications across physics and engineering.
His Orcid Number is 0000-0003-4858-7309

### v2

@@ V Concluding Remarks
Acknowledgements.
The Authors are grateful to an anonymous referee for a very careful reading of the manuscript and for the valuable suggestions to improve the whole presentation. L. R. E. gratefully acknowledges the hospitality of the Department of Molecular Sciences and Nanosystems at Ca’ Foscari University of Venice, Italy, and the Department of Materials, Environmental Sciences and Urban Planning (SIMAU) at the Polytechnic University of Marche (UnivPM), Ancona, Italy, where he served as a Visiting Scholar during the preparation of this work.
Appendix A Tangent to a Cycloid
The crucial point here is to consider the tangent BG to the cycloid as parallel to the chord AE of its generating circle, as depicted in Fig. 1, in order to relate the Huygens’s geometric derivation to Abel’s solution to the tautochrone problem. We start by defining
$y=AF$
as the height of point B. Thus, from the right triangle $\Delta DEA$, we may write
$AE=\sqrt{AF\dot{A}D}=\sqrt{2ry}$
and
$\sin\theta=\sin\angle EDA=\frac{2ry}{2r}=\sqrt{\frac{y}{2r}}.$
However, $\theta$ is also the angle the tangent BG makes with the horizontal, which allow us to write
$\sin\theta=\frac{dy}{ds}\quad{\textrm{and}}\quad\frac{dy}{ds}=\sqrt{\frac{y}{2r}}.$
This equation is essentially the same as Eq. 2, thus providing a link between the approaches of Abel and Huygens. In this framework, we can interpret Eq. 2 as stating that a necessary and sufficient condition for the curve $s=f(y)$ to be a tautochrone with quarter-period $T$ is that there exists a circle of radius
$r=\frac{gT^{2}}{\pi^{2}},$
whose tangents are parallel to the corresponding tangents of the tautochrone, having in mind Fig. 1.
Appendix B Why is the Cycloid a Tautochrone?
Let us try to illustrate in simple terms what makes the cycloid a tautochrone in a uniform gravitational field.
We consider a mass released from the rest and try to show that it oscillates in simple harmonic motion along the length of the cycloid. To do this, we recognize that arclength AB in Fig. 1 may be expressed as
$s=4r\sin\theta,$
where, as before, $\theta$ is the angle the tangent to the cycloid at B makes with the horizontal.
Now, the restoring force on a mass sliding along the cycloid (i.e., the component of its weight tangent to the curve), given by
$F=mg\sin\theta=\frac{mg}{4r}s,$
is proportional to its distance (which is measured along the cycloid) from equilibrium. This way, the mass oscillates in simple harmonic motion with a period
$T_{m}=2\pi\sqrt{\frac{4r}{g}}.$
Why this suggests that the cycloid is a tautochrone? The key point is that the arclength BA of the cycloid, as depicted in Fig. 1, is twice the length of the tangent segment BG. Since the latter, BG = $2r\sin\theta$ is proportional to $\sin\theta$, the same occurs with the former, which is precisely the condition for simple harmonic motion.
This indicates that the Abel’s proof says that the only way for a curve to be a tautochrone is that a mass released from the rest oscillates in a simple harmonic motion along the curve.
B.1 The Authors
Luiz Roberto Evangelista
is Professor of Theoretical Physics at the PhD Preogram in Physics at the University of Maringá (UEM), Brazil. His research interest lie in complex fluids, complex systems, and history of physics, and include mathematical physics of liquid-crystals, diffusion problems, and adsorption-desorption phenomena.
Francesco Mainardi
is Professor (Retired) of Mathematical Physics at the University of Bologna, renowned for his pioneering work in fractional calculus. His research focuses on fractional differential equations, viscoelasticity, and anomalous diffusion. He played a key role in promoting the Mittag-Leffler and Wright functions in applied contexts. Mainardi authored influential books, including Fractional Calculus and Waves in Linear Viscoelasticity. He remains an active voice in the field, shaping theory and applications across physics and engineering.
B.2 Financial Support
@@ B.4 Conflict of Interest
B.5 Data Availability Statement
@@
B.6 Generative Artificial Intelligence (AI)
This manuscript benefited from the use of Generative AI tools (ChatGPT and DeepSeek) for translation assistance and grammar refinement, given that the authors are not native speakers of English.